# Emotion Granularity from Text: An Aggregate-Level Indicator of Mental Health

### v1

@@ 4.2 Quantifying Emotion Granularity
    ADHD–control lower lower higher lower lower
@@
    MDD–control lower lower lower lower
@@
Table 2: Emotion Granularity - using Utterance-level Window: The difference in emotion granularity between each MHC group and the control. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference in granularity. E.g., ‘lower’ for ADHD–control under $EG(pos)$ means that the ADHD group has significantly lower positive emotion granularity than the control group.
5 Examining the Link between Emotion Granularity and Mental Health
After preprocessing the posts for each user and computing their emotion granularity scores for the various emotions groups, we performed statistical analyses to determine if there were significant differences between the emotion granularities of each of the MHC groups and the control group.
To test this, we performed an independent t-test between each MHC and the control. However, in order to perform this test the data must meet certain assumptions. These assumptions are explained further in Appendix B. In summary, the data is approximately normally distributed and the groups are independent of each other; however the data does not meet the equal variance (i.e., homoscedasticity) assumption (therefore we specify this when performing the t-test).
We performed independent t-tests to compare emotion granularities between each of the MHCs and the control group for each emotion group using the SciPy library Virtanen et al. (2020). To correct for multiple comparisons (eight tests performed for each MHC per emotion granularity group), we used the Benjamini–Hochberg procedure in the statsmodels library Seabold and Perktold (2010).
5.1 Term Specificity as a Control
Lower emotion granularity occurs when, for a person, the concepts of the relevant emotions are so broad (and non-specific) that their meanings overlap substantially. This work is testing the hypothesis of whether people who have self-disclosed as having an MHC have lower emotion granularity than those that do not. However, another plausible hypothesis is that people in a particular group (e.g., MHC or the control) tend to use more specific words overall. Doing so would imply a higher specificity (i.e, a higher granularity) in their usage of all words, and that the high granularity of emotion words is simply a by-product of their general style of speaking (posting online).
To ensure that the level of word specificity does not differ between MHCs and the control group and act as a confounder for our measure of emotion granularity, we performed a control experiment. We compute the average information content of the noun and verb terms in the posts of users in each group, and use this as a measure of the specificity of their language. The information content of terms has been studied in NLP for applications involving semantic similarity, particularly in information retrieval Resnik (1995); Sánchez et al. (2011).
We use the metric proposed in Resnik (1995), and implemented in the NLTK WordNet library, [^9] https://github.com/nltk/wordnet which combines information about the depth of the term in the WordNet tree hierarchy and its frequency of occurrence in a large corpus (here, the Brown corpus) to compute an information content score Miller (1995). We lemmatize each word using the inbuilt WordNet Morphy library, and consider the most frequent synset to obtain the information content score. We then compute the following measures of term specificity:
  • 6 Results
    In Table 2 we report the statistical results from the pairwise comparisons between each MHC and the control group, for the control experiment on general term specificity as well as emotion granularity, when scores are computed at the utterance-level. All statistically significant differences between an MHC and the control group are described as either ‘higher’ or ‘lower’, and a dash (–) for no statistical difference. A ‘lower’ value in a cell indicates that the MHC (rows) has lower emotion granularity (or lower term specificity) than the control group, i.e., higher correlation between emotion pairs in that group (columns); ‘higher’ indicates the MHC has higher emotion granularity (or higher term specificity) than the control group (i.e., lower correlation between emotion pairs in that group). Below we summarize the results for each column.
    IC(n) and IC(v): We do not see any significant differences in the information content of noun and verb terms ($IC(n)$ and $IC(v)$) between MHCs and the control group. This indicates that no group tends to use more specific or less specific language in general when posting on these platforms. Details on the statistical results are shown in Appendix E.
    EG(pos): All MHCs except for PPD had significantly lower positive emotion granularity than the control group (which had similar granularity compared to the control group). That is, these MHCs (ADHD, Anxiety, etc.) consistently expressed multiple positive emotions concurrently, more so than the control group.
    EG(neg): All MHCs had significantly lower negative emotion granularity than the control group, in both datasets. Thus these MHCs were generally not differentiating between the negative emotions of anger, disgust, fear, and sadness, as well as compared to the control group.
    EG(var): All MHCs except ADHD had significantly lower variable emotion granularity than the control group (i.e., these MHCs generally differentiated between surprise and anticipation less than the control group). On the other hand, ADHD had significantly higher variable emotion granularity than the control group (i.e., the ADHD group was better at differentiating between surprise and anticipation).
    EG(overall): All MHCs had significantly lower emotion granularity for emotion categories that express the same valence (the mixed valence emotions of surprise and anticipation are also included here). These MHCs are therefore expressing multiple close emotions frequently with one another – more so than the control group.
    EG(cross): ADHD, Anxiety, OCD, PTSD, and Depression (in both datasets) have significantly lower granularity between emotion pairs that come from different valence groups. This indicates that positive and negative emotions are expressed together more frequently in these MHCs compared to the control, as well as emotions like joy (positive valence) and surprise (variable valence).
    Aligning with self-report studies in psychology, the emotion granularity between negative-valence emotions is lower for all MHCs with utterance-level operationalizations. Positive emotion granularity is also correlated with many of the MHCs (except PPD). In general, the granularity of emotional expression between within-group emotion pairs is lower for all MHC groups compared to the control in both datasets. This is in line with both the theoretical and conceptual links established in the psychology literature on emotion granularity and mental health: the ability to better differentiate between emotion concepts that are close to one another leads to more adaptive health behaviour.
    6.1 Additional Window Sizes
    We examine how the measure of emotion granularity between MHCs and the control changes when we compute emotion scores with larger window sizes (multiple tweets/posts are included in computing the emotion score at a single time point).
    Many of the utterance-level outcomes are replicated for negative, positive, and overall emotion granularity with window size 100. Some measures are no longer significantly different between certain MHCs and the control (e.g., $EG(overall)$ for PPD). We also find that $EG(cross)$ is higher for certain MHCs (Anxiety, PPD, PTSD, Depression in Twitter-STMHD) when compared to the control, i.e, users in the control group are expressing negative and positive emotions together more frequently than those in the MHCs. When the window size is increased to 500 words, this reversal of trends for cross-valence emotions persists.
    With larger window sizes, we end up capturing emotions expressed by the individual over longer time spans (tweets posted over the span of several hours or days), rather than co-endorsement at the same time. We hypothesize that these effects of dialecticism, where the Control group has a higher co-occurrence of cross-valence group emotions, are capturing the extent to which users balance negative emotions with positive emotions (and vice versa), and how people bounce back from negative events, find the silver linings, etc. The consistent effects with 100 and 500-word windows, and for several MHCs, makes this a promising area for future work.
    All emotion granularity measures with window sized 100 and 500 are reported in Tables 3 and 4 in Appendix C.
    6.2 Individual Emotion Pairs
    • All eight MHCs in the Twitter-STMHD dataset had a lower granularity (a higher correlation) for anger–disgust and anger–sadness in the negative valence group.
    • Among cross-group emotion pairs, all eight MHCs had a lower emotion granularity (higher correlation) between anticipation–fear and anticipation–sadness.
    • Certain MHCs had a higher emotion granularity (lower correlation) between the following cross-group emotions: anger–joy co-occurs more in the control than in Bipolar, Depression, MDD, PPD, PTSD; however, it co-occurs less in control compared to ADHD.
    • Similarly, sadness–surprise co-occur more in the control when compared to Anxiety, Bipolar, Depression, and PPD; but less when compared to ADHD and PTSD.
    Detailed results for each of the emotion pairs and all MHCs are in Appendix D, Table 5.
    Discussion: While lower granularity among certain emotion pairs consistently function as indicators of all MHCs, we also see many instances where MHCs have a higher granularity between the emotions when compared to the control. These findings are of interest to researchers studying the links between how emotions are expressed in text, and how they are vary with with different MHCs.
    7 Conclusion
    In this work, we operationalized for the first time a computational measure of emotion granularity that can be derived from the textual utterances of individuals. We applied this measure to two social media datasets of posts from individuals who have self-disclosed as having an MHC. Our findings showed that our measure of emotion granularity is significantly lower for all the MHC groups when compared to a control group at an aggregate-level, for emotion pairs that express similar valence. Specifically, we found significantly lower values of negative and overall emotion granularity for all eight MHCs represented in our dataset. Similarly, we found all MHCs except for PPD had lower positive emotion granularity than compared to the control group. Our work makes an important contribution towards deriving aggregate-level indicators of emotional health from the large amounts of utterance data available on social media platforms.
    Limitations
    Our work uses the social media utterances of individuals to derive measures of emotional expression that, at an aggregate level, are found to correlate with multiple mental health conditions. While we use datasets that were compiled by other researchers in the field, we stress that they may not be representative of the general population. Our methods therefore cannot be directly applied to make inferences on other datasets without a careful experimental validation first. The datasets we study rely on self-disclosures made on social media platforms; it is possible that users report only one such MHC but are diagnosed with others, or that they misrepresent their diagnoses. Further, the users in the control groups may include those who have chosen to simply not self-disclose on these platforms. This can occur due to many reasons, like social desirability Latkin et al. (2017b) or impression management Tedeschi (2013)). Nonetheless, since we draw inferences at an aggregate level, the methods used can overcome some amount of noisy data.
    The emotion lexicons that we use are some of the largest that exist with wide coverage and large number of annotators (thousands of people as opposed to just a handful). However, no lexicon can cover the full range of linguistic and cultural diversity in emotion expression. The lexicons are largely restricted to words that are most commonly used in Standard American English and the capture emotion associations as judged by American native speakers of English. See Mohammad (2020) for a discussion of the limitations and best-practises in the use of emotion lexicons.
    Ethics Statement
    Individuals vary considerably in how, and how well, they express their internal emotional states using language. Our method of assessing the emotional states of users based on their utterances may miss several linguistic cues of emotion expression, and may not account for individual variation or the extent to which these emotions are expressed on social media. The emotionality of one’s language may also be conveying information about the emotions of the speaker, the listener, or something or someone else mentioned in the utterances. See further discussions of ethical considerations when using computational methods for affective science in Mohammad (2020, 2022).
    Appendix A Twitter-STMHD Dataset
    Appendix B Statistical Assumptions
    • Independence of observations: Since the text stream of utterances come from different people, we can assume these are independent observations. [^10] In reality individuals are largely influenced by one another as we see interact and engage with content from various communities, which can influence our emotional state and therefore utterances. However, for the purposes of our experiments since the utterances come from different people we can assume they are independent.
    Appendix C Emotion Granularity: Various Window Sizes

### v2

@@ 4.2 Quantifying Emotion Granularity
    ADHD–control lower lower lower lower lower
    Anxiety–control lower lower lower lower lower
    Bipolar–control lower lower lower lower lower
    MDD–control lower lower lower
    OCD–control lower lower lower lower lower
    PPD–control lower
    PTSD–control lower lower lower lower lower
    Depression–control lower lower lower lower lower
Reddit eRisk
    Depression–control lower lower lower lower
Table 2: The difference in emotion granularity between each MHC group and the control. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference in granularity.
5 Emotion Granularity and Mental Health
We now test if there are significant differences between the emotion granularities of each of the MHC groups and the control group, using t-tests. We first limit the users in each group by placing thresholds on (a) the number of user tweets with a valid emotion score (set to a minimum of 50), and (b) the number of unique lexicon terms used in their tweets (set to a minimum of 25). These thresholds ensure that we are drawing inferences based on users with valid emotion arcs, with sufficient lexicon coverage and temporal information.
We performed independent t-tests to compare emotion granularities between each of the MHCs and the control group, for each emotion group, using the SciPy library Virtanen et al. (2020). To correct for multiple comparisons (eight tests performed for each MHC per emotion granularity group), we used the Benjamini–Hochberg procedure in the statsmodels library Seabold and Perktold (2010). Further details on the data assumptions for t-tests are in Appendix C.
5.1 Term Specificity as a Control
Lower emotion granularity occurs when, for a person, the concepts of the relevant emotions are so broad (and non-specific) that their meanings overlap substantially. This work is testing the hypothesis of whether people who have self-disclosed as having an MHC have lower emotion granularity than those that do not. However, another plausible hypothesis is that people in a particular group (e.g., MHC or the control) tend to use more specific words overall. Doing so would imply a higher specificity (i.e, a higher granularity) in their usage of all words, and that the high granularity of emotion words is simply a by-product of their general style of speaking (or posting online).
To ensure that the level of word specificity does not differ between MHCs and the control group and act as a confounder for our measure of emotion granularity, we performed a control experiment. We compute the average information content of the noun and verb terms in the posts of users in each group, and use this as a measure of the specificity of their language. We use the metric proposed in Resnik (1995), and implemented in the NLTK WordNet library, [^9] https://github.com/nltk/wordnet which combines information about the depth of the term in the WordNet tree hierarchy and its frequency of occurrence in a large corpus (here, the Brown corpus) to compute an information content score Miller (1995). We then compute the following measures of term specificity for each user:
  • 6 Results
    In Table 2 we report the statistical results from the pairwise comparisons between each MHC and the control group, for the control experiment on general term specificity as well as emotion granularity, when scores are computed at the utterance-level.
    All statistically significant differences between an MHC and the control group are described as either ‘higher’ or ‘lower’, and a dash (–) for no statistical difference. A ‘lower’ value in a cell indicates that the MHC (rows) has lower emotion granularity (or lower term specificity) than the control group, i.e., higher correlation between emotion pairs in that group (columns); ‘higher’ indicates the MHC has higher emotion granularity (or higher term specificity) than the control group (i.e., lower correlation between emotion pairs in that group). In Table 12 in the Appendix, we also report the absolute Spearman correlation scores for each group. Below we summarize the results for each column.
    6.1 Emotion Granularity as an Indicator of MHCs
    IC(n) and IC(v): We do not see any significant differences in the information content of noun and verb terms ($IC(n)$ and $IC(v)$) between MHCs and the control group. This indicates that no group tends to use more specific or less specific language in general when posting on these platforms. Details on the statistical results are shown in Appendix H.
    EG(pos): All MHCs except for PPD had significantly lower positive emotion granularity than the control group (which had similar granularity compared to the control group). That is, tweeters in these MHC groups (ADHD, Anxiety, etc.) consistently expressed multiple positive emotions concurrently, more so than the control group.
    EG(neg): All MHCs except MDD had significantly lower negative emotion granularity than the control group, in both datasets. Thus, tweeters in these MHCs were generally not differentiating between the negative emotions of anger, disgust, fear, and sadness, as well as the control group.
    EG(var): Tweeters in the ADHD, Anxiety, Bipolar, OCD, PTSD, and Depression (Twitter-STMHD) had significantly lower variable emotion granularity than the control group (i.e., these groups generally differentiated between surprise and anticipation less than the control group).
    EG(overall): All MHCs except PPD had significantly lower emotion granularity for emotion categories that express the same valence (the mixed valence emotions of surprise and anticipation are also included here). Tweeters in these groups are therefore expressing multiple close emotions frequently with one another – more so than the control group.
    EG(cross): All MHCs except PPD had significantly lower granularity between emotion pairs that come from different valence groups. This indicates that positive and negative emotions are expressed together more frequently by tweeters in these groups compared to the control, as well as emotions like joy (positive valence) and surprise (variable valence).
    Aligning with self-report studies in psychology, the emotion granularity between negative-valence emotions is lower for most (7 out of 8) MHCs in our datasets with utterance-level operationalizations. Positive emotion granularity is also correlated with many of the MHCs (7 out of 8 disorders). In general, the granularity of emotional expression between within-group emotion pairs is lower for all MHC groups compared to the control in both datasets, except PPD. This is in line with both the theoretical and conceptual links established in the psychology literature on emotion granularity and mental health: the ability to better differentiate between emotion concepts that are close to one another leads to more adaptive health behaviour.
    Variation with hyperparameter choices: We observe only minor variations from the results reported in Table 2 when the hyperparameters described in Section 4.1 for emotion arc construction were changed – less than 10% of the cells differed in their values across all variations. We provide a more detailed report in Appendix F.
    6.2 Additional Window Sizes
    We also examined how the measures of differences in emotion granularity between MHCs and the control change when we compute emotion scores with larger window sizes.
    Many of the utterance-level outcomes are replicated for negative, positive, and overall emotion granularity with window sizes 100 and 500. Some measures are no longer significantly different between certain MHCs and the control. We also find that $EG(cross)$ is higher for certain MHCs (Anxiety, PPD, PTSD, Depression in Twitter-STMHD) when compared to the control, i.e, users in the control group are expressing negative and positive emotions together more frequently than those in the MHCs.
    With larger window sizes, we end up capturing emotions expressed by the individual over longer time spans (tweets posted over the span of several hours or days), rather than co-endorsement at the same time. We hypothesize that these effects of dialecticism, where the control group has a higher co-occurrence of cross-valence group emotions, are capturing the extent to which users balance negative emotions with positive emotions (and vice versa). The consistent effects with 100 and 500-word windows, and for several MHCs, makes this a promising area for future work. All emotion granularity measures with window sized 100 and 500 are reported in Appendix F.1.
    6.3 Individual Emotion Pairs
    • Seven out of the eight MHCs in the Twitter-STMHD dataset had a lower granularity (a higher correlation) for anger–disgust (except PPD) and anger–sadness (except MDD) in the negative valence group.
    • All eight MHCs had a lower emotion granularity (higher correlation) between multiple cross-group emotion pairs, notably those involving the mixed-valence emotions of anticipation and trust.
    • Contrary to trends, the Bipolar group had a higher emotion granularity (i.e, a lower correlation of emotion arcs) for the cross-group emotion pairs of anger–joy and fear–joy.
    Detailed results for each of the emotion pairs and all MHCs are in Appendix G, Table 10.
    Discussion: While lower granularity among certain emotion pairs consistently function as indicators of all MHCs, we also see a few instances where MHCs (specifically Bipolar disorder) have a higher granularity between the emotions when compared to the control. These findings are of interest to researchers studying the links between how emotions are expressed in text, and how they vary with different MHCs.
    7 Conclusion
    In this work, we operationalized for the first time a computational measure of emotion granularity that can be derived from the textual utterances of individuals. We applied this measure to two social media datasets of posts from individuals who have self-disclosed as having an MHC. Our findings showed that our measure of negative emotion granularity is significantly lower for 7 out of the 8 MHC groups under consideration when compared to a control group, at an aggregate-level. Also, all MHCs except for PPD had lower overall emotion granularity (and lower positive emotion granularity) compared to the control group. Our work makes an important contribution towards deriving aggregate-level indicators of emotional health from the large amounts of utterance data available on social media platforms. We hope this opens up an avenue of future work to explore emotional expression in text and mental health.
    Limitations
    Our work uses the social media utterances of individuals to derive measures of emotional expression that, at an aggregate level, are found to correlate with multiple mental health conditions. While we use datasets that were compiled by other researchers in the field, we stress that they may not be representative of the general population. Our methods therefore cannot be directly applied to make inferences on other datasets without a careful experimental validation first. The datasets we study rely on self-disclosures made on social media platforms; it is possible that users report only one such MHC but are diagnosed with others, or that they misrepresent their diagnoses. Further, the users in the control groups may include those who have chosen to simply not self-disclose on these platforms. This can occur due to many reasons, like social desirability Latkin et al. (2017) or impression management Tedeschi (2013). Nonetheless, since we draw inferences at an aggregate level, the methods used can overcome some amount of noisy data.
    The emotion lexicons that we use are some of the largest that exist with wide coverage and large number of annotators (thousands of people as opposed to just a handful). However, no lexicon can cover the full range of linguistic and cultural diversity in emotion expression. The lexicons are largely restricted to words that are most commonly used in Standard American English and they capture emotion associations as judged by American native speakers of English. See Mohammad (2023) for a discussion of the limitations and best-practises in the use of emotion lexicons.
    Lastly, further work should explore if the relationships we found hold around various social factors such as age, region, language, etc. As we focus on English text, and the region of users is not known (some information could be extracted from user profiles in the Twitter-STMHD dataset however it is fairly noisy), conclusions should be drawn cautiously across various sociolinguistic factors.
    Ethics Statement
    Individuals vary considerably in how, and how well, they express their internal emotional states using language. Our method of assessing the emotional states of users based on their utterances may miss several linguistic cues of emotion expression, and may not account for individual variation or the extent to which these emotions are expressed on social media. The emotionality of one’s language may also be conveying information about the emotions of the speaker, the listener, or something or someone else mentioned in the utterances. See further discussions of ethical considerations when using computational methods for affective science in Mohammad (2023, 2022).
    Appendix A Twitter-STMHD Dataset
    Appendix B Lexicon Words Removed
    We considered the following sets of terms to be stop-words, which do not contribute to the emotion score of an utterance, for our analysis:
    • Common stopwords: We remove common English stopwords, such as the, of, for, etc. We use the list of English stopwords from the Python NLTK library. The full list can be found at https://gist.github.com/sebleier/554280.
    • Domain-specific stopwords: We remove terms (words and word pairs) whose dominant usage on social media platforms differs from their annotated sense (e.g, like, chaotic evil, good morning). The full list of these terms is in Table 3.
    • MHC-associated terms: Finally, we filter out terms that are explicitly associated with the MHCs that we consider, such as anxiety, mental health, and panic attack. The full list of terms is in Table 4.
    love flu shot raptor discord
    christmas good day good morning good evening
    birthday good night good afternoon bloody murder
    pretty true crime full time gut punch
    vibe wholesome content slur word life time
    vote jump scares hot chocolate chaotic evil
    trump fever dream chaotic energy chaotic good
    like guilty pleasure chaotic neutral hot mess
    Table 3: Twitter-specific words and bigrams removed from the emotion lexicons.
    disability ptsd
    psychosis adhd
    suicide depressive
    depressed disorder
    anxiety mental health
    anxious mental illness
    disabled panic attack
    Table 4: Mental health specific terms removed from the emotion lexicons.
    Appendix C Statistical Assumptions
    • Independence of observations: Since the text stream of utterances come from different people, we can assume these are independent observations. [^10] In reality individuals are largely influenced by one another as we see, interact and engage with content from various communities, which can influence our emotional state and therefore utterances. However, for the purposes of our experiments since the utterances come from different people we can assume they are independent.
    Appendix D Emotion Lexicons
    In Table 5, we report statistics on the number of emotion terms in each lexicon for the eight emotions we consider in this work, and the number of terms common to and mutually-exclusive between each emotion pair.
    emo1 emo2 e1-all e2-all e12-comm e1-excl e2-excl
    anger anticipation 1157 782 43 1114 739
    anger disgust 1157 886 407 750 479
    anger fear 1157 1343 551 606 792
    anger joy 1157 946 3 1154 943
    anger sadness 1157 1014 382 775 632
    anger surprise 1157 454 102 1055 352
    anger trust 1157 1332 6 1151 1326
    anticipation disgust 782 886 19 763 867
    anticipation fear 782 1343 82 700 1261
    anticipation joy 782 946 283 499 663
    anticipation sadness 782 1014 32 750 982
    anticipation surprise 782 454 131 651 323
    anticipation trust 782 1332 283 499 1049
    disgust fear 886 1343 400 486 943
    disgust joy 886 946 1 885 945
    disgust sadness 886 1014 336 550 678
    disgust surprise 886 454 56 830 398
    disgust trust 886 1332 2 884 1330
    fear joy 1343 946 2 1341 944
    fear sadness 1343 1014 545 798 469
    fear surprise 1343 454 137 1206 317
    fear trust 1343 1332 9 1334 1323
    joy sadness 946 1014 0 946 1014
    joy surprise 946 454 113 833 341
    joy trust 946 1332 308 638 1024
    sadness surprise 1014 454 73 941 381
    sadness trust 1014 1332 3 1011 1329
    surprise trust 454 1332 56 398 1276
    Table 5: Emotion Lexicons: For each emotion pair (emo1, emo2), the number of terms in each lexicon (e1-all, e2-all), the number of emotion terms common to the two lexicons (e12-comm), and the number of mutually-exclusive emotion terms (e1-excl, e2-excl).
    Appendix E Visualization of Emotion Arcs
    In Figure 1, we plot the emotion arcs for the anger-fear emotion pair, from the negative valence group, for a tweeter from an MHC group of the Twitter-STMHD dataset. Emotion scores are computed and plotted at the utterance-level, i.e, independently for each tweet by the user. Note that larger window sizes and overlapping windows will lead to smoother arcs.
    Figure 1: Emotion arcs: Tweet-level emotion arcs for anger and fear, for a sampled user from the Twitter-STMHD dataset.
    Appendix F Emotion Granularity: Hyperparameters
    We report the results of the statistical analyses of emotion granularity when emotion arcs were generated using different choices of the hyperparameters described in Section 4.1. Table 6 reports the results when non-lexicon terms (and tweets) are assigned a score of 0, and only mutually-exclusive emotion terms are considered, similar to Table 2, but no user thresholds on number of tweets and unique emotion terms are applied. Table 7 reports the results when non-lexicon terms (and tweets) are not considered, and user thresholds are set to 50 and 25, (similar to Table 2), and only mutually-exclusive emotion terms are considered (similar to Table 2).
    We find that largely the results do not change. However, when non-lexicon terms and tweets are ignored, this results in a smaller set of tweets to compute the emotion arc over, and fewer tweeters who meet the user thresholds for each group. This results in signals turning off for certain MHCs.
    Twitter-STMHD
        ADHD–control lower lower lower lower lower
@@ Appendix F Emotion Granularity: Hyperparameters
        MDD–control lower lower lower
@@
    Table 6: Emotion Granularity - hyperparameter variations: The difference in emotion granularity between each MHC group and the control. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference in granularity. Non-lexicon terms and tweets are assigned a score of zero; user tweet and unique term thresholds are both set to 0, and only mutually-exclusive emotion terms are considered.
    Twitter-STMHD
        ADHD–control lower lower lower lower lower
        Anxiety–control lower lower lower lower
        Bipolar–control lower lower lower lower lower
        MDD–control
        OCD–control lower lower lower lower lower
        PPD–control lower lower lower lower lower
        PTSD–control lower lower lower lower lower
        Depression–control higher lower lower lower
    Reddit eRisk
        Depression–control lower lower lower lower
    Table 7: Emotion Granularity - hyperparameter variations: The difference in emotion granularity between each MHC group and the control. A significant difference is indicated by the word ‘lower’ or ‘higher’, indicating the direction of the difference in granularity. Non-lexicon terms and tweets are discarded; user tweet and unique term thresholds are set to 50 and 25, and only mutually-exclusive emotion terms are considered.
    F.1 Various Window Sizes
    We report the results of the statistical analyses of emotion granularity when emotion arcs were generated using two other window sizes: 100 (Table 8) and 500 (Table 9). All other hyperparameters are the same as for Table 2.